# Breaking the Curse of Dimensionality: On the Stability of Modern Vector Retrieval

### v1

@@ 1 Introduction
The success of modern vector retrieval raises a question that we find surprisingly unexplored in the current research literature: Given that vector databases efficiently search high-dimensional embedding spaces, what properties do these search algorithms and embeddings possess that make them amenable to sub-linear time search algorithms? In addition to enhancing our theoretical understanding of the curse of dimensionality, successfully answering this question also has implications for machine learning model developers and vector database engineers. In particular, if we can identify necessary or sufficient conditions for embedding-based search to “break“ the curse of dimensionality, practitioners can potentially leverage these insights when working to optimize models and search algorithms by maintaining the essential invariants that account for the success of these representations in high dimensions.
@@
    Multi-Vector Search which involves performing near-neighbor search in the case where both the query and the database items consist of sets of vectors. This style of search has become extremely popular due to the recent success of “late interaction” models such as ColBERT (Khattab & Zaharia, 2020).
@@
    Filtered Vector Search which involves near neighbor search in the presence of discrete attribute filters. This setting has also become widely deployed in vector databases in practice (Gollapudi et al., 2023; Wang et al., 2023) due to the need to filter results based on discrete metadata.
@@
    Sparse Vector Search which is near neighbor search in the special case where the embedding vectors are extremely high-dimensional (such as in the tens or hundreds of thousands) but only a small number of components are non-zero. This search style has become popular due to the success of learned sparse representation models such as SPLADE (Formal et al., 2021a; b) and semantic models in recommender systems (Rajput et al., 2023; Penha et al., 2025).
Intuitively, stability captures the notion that a slight perturbation to a query should not dramatically change the list of near neighbors to that query. The landmark paper of (Beyer et al., 1999) and a subsequent follow-up work of (Durrant & Kabán, 2009) together establish a necessary and sufficient condition for stability via a quantity called relative variance. However, to our knowledge, no work has studied extending these prior results to modern vector database settings, particularly in the three cases that we focus on in this paper. Thus, our work fills a gap in the literature in understanding why modern neural representations enable efficient sub-linear vector search in high dimensions. Our specific contributions are as follows:
  • We prove that multi-vector search with the commonly-used Chamfer distance is stable if the induced single vector search problem is strongly stable (along with some technical conditions on the composition of the vector sets). However, not all aggregations are stability preserving; we give a counter-example showing that stability does not hold in general for average pooling. This result provides a theoretical justification for the choice of the Chamfer distance used in the popular ColBERT family of models (Khattab & Zaharia, 2020; Santhanam et al., 2022; Faysse et al., ).
@@
  • In addition to our theoretical results, we also empirically validate each of these theorems on a combination of synthetic and real embedding datasets.
@@ 2 Background
In this section, we will introduce the necessary mathematical preliminaries for our main results. We begin with a formal definition of nearest-neighbor search.
@@
A near-neighbor search problem is a 3-tuple $(Q,D,\delta)$ for a query set $Q$ and database collection $D$ where $q,d\in\mathbb{R}^{m}$ for all $q\in Q$ and $d\in D$. For each $q\in Q$, we use the distance metric $\delta$ to compute
@@
In the remainder of this paper, we will assume that all database elements $d\in D$ are sampled from some distribution and all query elements $q\in Q$ are sampled independently from some distribution. Under this probabilistic model, the distance $\delta(q,d)$ for $q\sim Q$ and $d\sim D$ is a random variable. In a seminal paper, Beyer et al. (1999) introduced this probabilistic model to provide a precise formalization of the curse of dimensionality in the context of near-neighbor search by defining the notion of instability.
@@
Example 2.5 (Example of Stability).
One concrete example of a stable near-neighbor search problem, originally proposed by (Beyer et al., 1999), arises from constructing query and document vectors such that each component in every vector is equal, and this value is iid sampled from some nontrivial distribution.
Example 2.6 (Example of Instability).
An example of an unstable search problem, also originally identified by (Beyer et al., 1999), is one where the query and document distributions have finite moments and each component of all vectors is iid sampled from these respective distributions.
To prove that a near-neighbor problem is stable or unstable, we will also stand on the shoulders of two prior theorems from (Beyer et al., 1999) and (Durrant & Kabán, 2009), respectively.
@@
Remark 2.9.
We can equivalently state the condition in Theorem 2.8 as $\lim_{m\to\infty}\frac{\mathrm{Var}\left[\delta(q,d)\right]}{\mathbb{E}\left[\delta(q,d)\right]^{2}}>0$ since the relative variance is non-negative by definition. We will use both characterizations interchangeably in this work.
@@ 3.1 Conditions for Stability
The paper of (Beyer et al., 1999) was a foundational work that inspired significant follow-up research in understanding the conditions for stable vector search. (Aggarwal et al., 2001) study stability under common $\ell_{p}$ norm distance metrics, where they find that $\ell_{1}$ and fractional distance metrics are theoretically and practically favorable to $\ell_{2}$ distances. Furthermore, (Houle et al., 2010) prove that the so-called shared-neighbor distance function, which measures the number of near-neighbors in common between two points, is a more stable measure than absolute distance since the latter loses discriminatory power in high dimensions. In a different vein, (Hui & Gao, 2021) study stability in the context of sequential data, such as genomes and sensor readings, and show that, under uniformly distributed sequences, the RelVar under edit distance goes to zero in the limit. Moreover, (Pestov, 2013) extends the results of (Beyer et al., 1999) from $k$-nearest neighbor search to $k$-nearest neighbor classification where they find that instability of vector search also implies barriers to the learnability of near-neighbor classification.
Our work is a direct follow-up to this rich prior literature in understanding the conditions for stable vector search. We focus on settings that have recently gained significant popularity amongst practitioners and provide an explanation for why these high-dimensional search settings succeed in spite of the curse of dimensionality.
3.2 Intrinsic Dimensionality
A closely related concept to stability is intrinsic dimensionality (ID) which measures the dimensionality of the subspace that a collection of vectors lie within. Intuitively, if a database of high-dimensional vectors lies within a much lower-dimensional manifold, then a near-neighbor search over these points is insulated from the curse of dimensionality effect. Thus, much like stability, ID provides a foundational theoretical tool for explaining why modern vector retrieval is capable of circumventing the curse of dimensionality. Consequently, ID has been well-studied with multiple proposed formal definitions. Perhaps the most influential definition of ID is that of Karger & Ruhl (2002) via a geomeric property called the expansion rate. Another popular definition heavily utilized in near-neighbor search algorithm design is the notion of doubling dimension proposed by Gupta et al. (2003). We refer interested readers to the recent book of Bruch (2024) for the precise definitions. More recently, several works have measured the intrinsic dimensionality of modern neural embeddings and have reported that the ID is often much less than the ambient dimensionality. For instance, Razzhigaev et al. (2024) study the intrinsic dimensionality of embeddings from transformer models (Vaswani et al., 2017) and find that the ID typically lies between 20 and 60 over the course of pretraining. In addition, Gong et al. (2019) compute the intrinsic dimensionality of popular image model embeddings, finding that the ID in practice typically ranges from 12 to 19, well below the ambient dimensions of 128 to 512. Moreover, Ansuini et al. (2019) go beyond embeddings and measure the ID of convolutional neural network (CNN) layers where they find the ID is an order of magnitude smaller than the ambient dimension.
Given that neural models consistently exhibit low intrinsic dimensionality, a natural question to ask is whether we still need to study why modern vector databases succeed in spite of the curse of dimensionality. On the contrary, we argue that intrinsic dimensionality alone is insufficient to explain the phenomena we see in practice. First and foremost, as we empirically measure in Section 4, a dimensionality as low as 32, which is within the empirical ID measurements of real embeddings, is still high enough for the curse of dimensionality to diminish the search quality of approximate near neighbor search algorithms. This observation motivates the need to study stability explicitly. Secondly, modern embeddings produced by large language models and used in practice continue to grow in dimension into the thousands and tens of thousands of components. It is not clear if the intrinsic dimensionality will remain low as these embedding models scale and yet we observe they remain amenable to sub-linear vector search techniques. Moreover, stability analysis provides actionable guidance that ID alone cannot. For instance, our multi-vector stability theorem explains why ColBERT’s choice of Chamfer distance succeeds where average pooling might fail, a distinction invisible to ID measurements which only characterize the embedding space geometry, not the aggregation function. Similarly, our filtered search results suggest that practitioners can induce stability through sufficiently large penalties, even when the underlying embeddings are intrinsically high-dimensional.
@@ 3.3 Multi-Vector Search
The first retrieval setting we consider in this work is multi-vector search, where each query and document may be represented as sets of vectors under a chosen set-similarity function. This style of vector retrieval first gained traction with the celebrated ColBERT paper (Khattab & Zaharia, 2020). Since the introduction of ColBERT, multiple works have studied how to accelerate of multi-vector search algorithms, including PLAID (Santhanam et al., 2022), DESSERT (Engels et al., 2023), and MUVERA (Jayaram et al., 2024). In addition, Gollapudi et al. (2025) demonstrate that multi-vector search under the commonly-used Chamfer distance (Barrow et al., 1977) can be integrated into the existing graph-based DiskANN index (Jayaram Subramanya et al., 2019) originally conceived for single-vector search. Furthermore, Faysse et al. extend the multi-vector modeling paradigm pioneered by ColBERT beyond text to multimodal vision-language retrieval. However, despite the tremendous research interest in multi-vector search, we note that no prior work has studied how the use of vector sets relates to the curse of dimensionality. We address this gap in the literature.
@@ 3.4 Filtered Similarity Search
With the advent of vector databases in retrieval applications, there has also been a surge of interest in augmenting traditional near-neighbor seach indexes with the capability of filtering points according to additional metadata. For instance, an e-commerce search service might want to return relevant near-neighbor results subject to additional attributes associated to each point such as “free shipping eligible.” Consequently, a number of works have proposed techniques for augmenting traditional vector search algorithms with filtering capabilities, such as NHQ (Wang et al., 2023), Airship (Zhao et al., 2022), Filtered DiskANN (Gollapudi et al., 2023), Acorn (Patel et al., 2024), CAPS (Gupta et al., 2023), IVF2 Landrum et al. (2025), and approximate window search (Engels et al., 2024).
@@ 3.5 Sparse Embedding Search
We conclude this section by turning our attention to a third form of embedding-based retrieval that has recently gained significant interest, namely sparse vector search. A sparse vector is one which is typically extremely high-dimensional (at least tens of thousands of components) but with only a small number of non-zero entries. In fact, sparse vectors predate the recent advent of neural representation learning as traditional keyword retrieval via inverted indexes (Schütze et al., 2008) can also be viewed as sparse vector search. Recently, sparse vectors have seen a renewal of interest in the form of learned sparse embeddings as pioneered by the influential SPLADE (Formal et al., 2021b; a) family of models that combine the generalization power of neural representations with the interpretability properties of sparse vectors. Introduced in 2024, the Seismic inverted index-based approach (Bruch et al., 2024) remains the current state-of-the-art for efficient and scalable sparse vector search. The key insight driving the design of Seismic is a novel observation that learned sparse embeddings, such as SPLADE, exhibit a property called concentration of importance, which states that the total $\ell_{1}$ mass of the sparse vector is primarily determined by a small subset of the non-zero coordinates. While the Seismic authors leverage this new concentration of importance observation to design an efficient indexing scheme, they do not consider the question of why learned sparse embedding search succeeds in spite of the extremely high vector dimensionalities. We take up that question in this paper and, inspired by this prior work, formalize the notion of concentration of importance. Although it turns out to not be sufficient for sparse vector search stability, we show that, under the additional assumption of overlap of importance and some mild technical assumptions, we can prove stability.
4 Practical Implications of Stability
In this section, we empirically demonstrate that stability is not merely a theoretical curiosity but also has substantial bearing on the quality of vector retrieval algorithms in practice. To our knowledge, this is the first empirical demonstration of the effects of instability on popular approximate ANN search algoritms. In Figure 1, we report the results of running two of the most popular ANN search techniques, namely the hierarchical navigable small worlds (HNSW) graph-based index (Malkov & Yashunin, 2018) and the inverted file (IVF) partitioning-based index (Douze et al., 2024), across stable and unstable datasets.
We fix a standard configuration of hyperparameters for each respective algorithm and compute the top-10 recall by taking the average of 1000 query vectors over a database of 1 million document vectors. To generate the unstable datasets at varying dimensionalities, we sample the value of each component independently from a univariate standard Gaussian distribution following the classic construction of (Beyer et al., 1999). To construct the stable datasets, we generate vectors such that they belong to one of five distinct clusters. [^1] Our experimental code is available at https://github.com/vihan-lakshman/ann-stability-theory. For consistency, we fix the search index hyperparameters across all experiments; for HNSW, we set ef-construction and ef-search to be $200$ and $\texttt{M}=16$; for IVF, we set the number of clusters to $\texttt{nlist}=\sqrt{n}$, where $n$ is the number of database elements and the number of probes to $\texttt{nprobes}=\frac{\sqrt{n}}{4}$.
From Figure 1, we see that these approximate near-neighbor search algorithms severely suffer in recall over unstable databases since the lack of contrast between points renders sublinear algorithms that avoid a brute force scan ineffective. We also see that the impact of instability worsens in higher dimensions. These results underscore the importance of stability in practice. We also observe that IVF is more robust to instability than HNSW since IVF performs a full linear scan within the probed partitions. However, in practice, this style of search comes at an increase in latency compared to graph-based methods like HNSW.
@@ 5 Stability of Multi-vector Search
We now turn our attention to the stability of multi-vector search. We begin by introducing a formal mathematical model of multi-vector search including a new, analogous definition of stability in this setting. We then proceed to our main theorem of this section identifying a sufficient condition for such stability to exist. In summary, we make the following contributions in this section:
  1. 1.
    We introduce formal definitions for multi-vector search and multi-vector stability.
  2. 2.
    We generalize the theorems of (Beyer et al., 1999) and (Durrant & Kabán, 2009) and prove that a multi-vector search instance is stable if and only if the relative variance over the associated set aggregation distance function is positive.
  3. 3.
    We prove that multi-vector search with Chamfer distance is stable provided that the induced single-vector search problem (a near neighbor search instance over the singleton vectors) is stable. We further give a counterexample showing multi-vector search with average pooling may not be stable even if the induced single-vector search instance is stable. Finally, we also validate our results experimentally.
@@ 5.1 Theoretical Model
With this definition in hand, we can now state our first theorem of this section generalizing the results of (Beyer et al., 1999) and (Durrant & Kabán, 2009) to the multi-vector setting.
@@
A multi-vector search instance is stable if and only if
@@
Proof sketch.
This result follows directly from the original arguments of Beyer et al. (1999) and Durrant & Kabán (2009). In particular, the original proofs only need to assume that the distance metric is a non-negative scalar random variable. Since, by definition, a valid set aggregation function also meets these criteria, the result is immediate. For completeness, we provide a full proof in Appendix A which is a synthesis of the proofs presented in the original papers of Beyer et al. (1999) and Durrant & Kabán (2009). ∎
With this extension of stability to the multi-vector setting in hand, we will now define a procedure for constructing a single-vector search instance from a multi-vector problem.
@@ 5.2 Theoretical Results
A near-neighbor search problem $(Q,D,\delta)$ is $c$-strongly stable for some constant $c>1$ if $\frac{\mathrm{DMAX}}{\mathrm{DMIN}}>c$ almost surely.
Intuitively, strong stability asserts that a near neighbor problem is not only stable but also stable with a gap. We will need this stronger property when analyzing the stability of multi-vector search because, even if the induced single-vector search problem is stable, the organization of document vectors into sets might contract the maximum distance gap seen in the single vector search problem. Strong stability ensures that, as long as the vector sets obey a certain mild non-degeneracy condition, we retain stability even if the stability gap contracts.
@@
We provide the proofs of Lemma 5.8 and Theorem 5.9 in Appendix A.
5.3 Empirical Validation of Theorem Assumptions
Theorem 5.9 relies on three assumptions to prove the stability of multi-vector search under Chamfer distance: $c$-strong stability, non-degenerate document sets, and a non-negative bound on the sum of the covariances of the individual Chamfer distance terms. In this section, we experimentally demonstrate that all three of these assumptions are satisfied in practice, which underscores the practical utility of our result. Specifically, we use ColBERT embeddings Khattab & Zaharia (2020) to construct query and document vector sets for five standard English-based information retrieval datasets from the BEIR benchmark Thakur et al. . For each dataset, we empirically estimate the strong stability constant $c$ as the smallest value of the ratio between the maximum distance and the minimum distance across all queries in the induced single-vector search instance. We then use this value of $c$ to check the proportion of document sets that satisfy the non-degeneracy condition. Finally, we then compute the sum of the covariances of the near-neighbor distances of individual query vectors in a given query set and verify that the sum is non-negative. This non-negative covariance assumption is natural in practice since the tokens in a query set tend to be topically related. We summarize our empirical measurements in Table 1 where we see that all three theorem conditions are comfortably satisfied in practice.
Dataset
Stability
Constant ($c$)
Non-Degeneracy
Pass Rate (%)
Cov. Bound
Pass Rate (%)
Natural Questions 2.36 100.0 100.0
HotpotQA 2.12 100.0 100.0
TREC-COVID 2.59 100.0 100.0
NFCorpus 1.91 100.0 100.0
Requirement >1>1 =100=100 =100=100
5.4 Stability of ColBERT Embeddings

### v2

@@ 1 Introduction
The success of modern vector retrieval raises a question that we find surprisingly unexplored in the current research literature: Given that vector databases efficiently search high-dimensional embedding spaces, what properties do these search algorithms and embeddings possess that make them amenable to sub-linear time search algorithms? In addition to enhancing our theoretical understanding of the curse of dimensionality, successfully answering this question also has implications for machine learning model developers and vector database engineers. In particular, if we can identify necessary or sufficient conditions for embedding-based search to “break” the curse of dimensionality, practitioners can potentially leverage these insights when working to optimize models and search algorithms by maintaining the essential invariants that account for the success of these representations in high dimensions.
@@
    Multi-Vector Search which involves performing near-neighbor search in the case where both the query and the database items consist of sets of vectors. This style of search has become extremely popular due to the recent success of “late interaction” models such as ColBERT (Khattab and Zaharia, 2020).
@@
    Filtered Vector Search which involves near neighbor search in the presence of attribute filters. This setting has also become widely deployed in vector databases in practice (Gollapudi et al., 2023; Wang et al., 2023) due to the need to filter results based on external metadata.
@@
    Sparse Vector Search which is near neighbor search in the special case where the embedding vectors are extremely high-dimensional (such as in the tens or hundreds of thousands) but only a small number of components are non-zero. This search style has become popular due to the success of learned sparse representation models such as SPLADE (Formal et al., 2021a, b) and semantic models in recommender systems (Rajput et al., 2023; Penha et al., 2025).
Intuitively, stability captures the notion that a slight perturbation to a query should not dramatically change the list of near neighbors to that query. The landmark paper of (Beyer et al., 1999) and a subsequent follow-up work of (Durrant and Kabán, 2009) together establish a necessary and sufficient condition for stability via a quantity called relative variance. However, to our knowledge, no work has studied extending these prior results to modern vector database settings, particularly in the three cases that we focus on in this paper. Thus, our work fills a gap in the literature in understanding why modern neural representations enable efficient sub-linear vector search in high dimensions. Our specific contributions are as follows:
  • We prove that multi-vector search with the commonly-used Chamfer distance is stable if the induced single vector search problem is strongly stable (along with some technical conditions on the composition of the vector sets). However, not all aggregations are stability preserving; we give a counter-example showing that stability does not hold in general for average pooling. This result provides a theoretical justification for the choice of the Chamfer distance used in the popular ColBERT family of models (Khattab and Zaharia, 2020; Santhanam et al., 2022; Faysse et al., 2025).
@@
  • In addition to our theoretical results, we also empirically validate each of these theorems on a combination of synthetic and real embedding datasets [^1] https://github.com/vihan-lakshman/ann-stability-theory.
1.1 Additional Results in Appendix
We defer several results to the appendix. In Appendix A, we discuss prior work on intrinsic dimensionality, a related concept to the theory of stability. In Appendix B, we demonstrate empirically that stability is not merely a theoretical curiosity but rather has substantial bearing on the quality of vector databases in practice. In the remaining Appendices (C, D, and E), we provide full proofs for all of the new mathematical results introduced in this paper, and we empirically verify that the conditions of our theorems hold on real embedding datasets.
@@ 2 Background
In this section, we introduce the necessary mathematical preliminaries for our main results. We begin with a formal definition of nearest-neighbor search.
@@
A near-neighbor search problem is a 3-tuple $(Q,D,\delta)$ for a query set $Q$ and database collection $D$ where $q,d\in\mathbb{R}^{m}$ for all $q\in Q$ and $d\in D$. For each $q\in Q$, we use the distance metric $\delta\colon\mathbb{R}^{m}\times\mathbb{R}^{m}\to\mathbb{R}_{\geq 0}$ to compute
@@
In the remainder of this paper, we will assume that all database elements $d\in D$ and all query elements $q\in Q$ are sampled from two respective distributions. Under this probabilistic model, the distance $\delta(q,d)$ for $q\sim Q$ and $d\sim D$ is a random variable. In a seminal paper, Beyer et al. (1999) introduced this probabilistic model to provide a precise formalization of the curse of dimensionality in the context of near-neighbor search by defining the notion of instability.
@@
To prove that a near-neighbor problem is stable or unstable, we will also stand on the shoulders of two prior theorems from (Beyer et al., 1999) and (Durrant and Kabán, 2009), respectively.
@@ 3.1 Conditions for Stability
The paper of Beyer et al. (1999) was a foundational work that inspired significant follow-up research in understanding the conditions for stable vector search. Aggarwal et al. (2001) study stability under common $\ell_{p}$ norm distance metrics, where they find that $\ell_{1}$ and fractional distance metrics are theoretically and practically favorable to $\ell_{2}$ distances. Furthermore, Houle et al. (2010) prove that the so-called shared-neighbor distance function, which measures the number of near-neighbors in common between two points, is a more stable measure than absolute distance since the latter loses discriminatory power in high dimensions. In a different vein, Hui and Gao (2021) study stability in the context of sequential data, such as genomes and sensor readings, and show that, under uniformly distributed sequences, the RelVar under edit distance goes to zero in the limit. Moreover, Pestov (2013) extends the results of Beyer et al. (1999) from $k$-nearest neighbor search to $k$-nearest neighbor classification where they find that instability of vector search also implies barriers to the learnability of near-neighbor classification.
@@ 3.2 Multi-Vector Search
The first retrieval setting we consider in this work is multi-vector search, where each query and document may be represented as sets of vectors under a chosen set-similarity function. This style of vector retrieval first gained traction with the celebrated ColBERT paper (Khattab and Zaharia, 2020). Since the introduction of ColBERT, multiple works have studied how to accelerate of multi-vector search algorithms, including PLAID (Santhanam et al., 2022), DESSERT (Engels et al., 2023), and MUVERA (Jayaram et al., 2024). In addition, (Gollapudi et al., 2025) demonstrate that multi-vector search under the commonly-used Chamfer distance (Barrow et al., 1977) can be integrated into the existing graph-based DiskANN index (Jayaram Subramanya et al., 2019) originally conceived for single-vector search. Furthermore, Faysse et al. (2025) extend the multi-vector modeling paradigm pioneered by ColBERT beyond text to multimodal vision-language retrieval. However, despite the tremendous research interest in multi-vector search, we note that no prior work has studied how the use of vector sets relates to the curse of dimensionality. We address this gap in the literature.
@@ 3.3 Filtered Similarity Search
With the advent of vector databases in retrieval applications, there has also been a surge of interest in augmenting traditional near-neighbor seach indexes with the capability of filtering points according to additional metadata. For instance, an e-commerce search service might want to return relevant near-neighbor results subject to additional attributes associated to each point such as “free shipping eligible.” Consequently, a number of works have proposed techniques for augmenting traditional vector search algorithms with filtering capabilities, such as NHQ (Wang et al., 2023), Airship (Zhao et al., 2022), Filtered DiskANN (Gollapudi et al., 2023), Acorn (Patel et al., 2024), CAPS (Gupta et al., 2023), IVF2 (Landrum et al., 2025), and approximate window search (Engels et al., 2024).
@@ 3.4 Sparse Embedding Search
We conclude this section by turning our attention to a third form of embedding-based retrieval that has recently gained significant interest, namely sparse vector search. A sparse vector is one which is typically extremely high-dimensional (at least tens of thousands of components) but with only a small number of non-zero entries. In fact, sparse vectors predate the recent advent of neural representation learning as traditional keyword retrieval via inverted indexes (Schütze et al., 2008) can also be viewed as sparse vector search. Recently, sparse vectors have seen a renewal of interest in the form of learned sparse embeddings as pioneered by the influential SPLADE (Formal et al., 2021b, a) family of models that combine the generalization power of neural representations with the interpretability properties of sparse vectors. Introduced in 2024, the Seismic inverted index-based approach (Bruch et al., 2024) remains the current state-of-the-art for efficient and scalable sparse vector search. The key insight driving the design of Seismic is a novel observation that learned sparse embeddings, such as SPLADE, exhibit a property called concentration of importance, which states that the total $\ell_{1}$ mass of the sparse vector is primarily determined by a small subset of the non-zero coordinates. Inspired by this prior work, we formalize the notion of concentration of importance in this paper and draw a connection between this concept and stability. Although it turns out to not be sufficient for sparse vector search stability, we show that, under the additional assumption of overlap of importance and some mild technical assumptions, we can prove stability.
@@ 4 Stability of Multi-vector Search
We now turn our attention to the stability of multi-vector search. We begin by introducing a formal mathematical model of multi-vector search including a new, analogous definition of stability in this setting. We then proceed to our main theorem of this section identifying a sufficient condition for such stability to exist.
@@ 4.1 Theoretical Model
We can now state our first theorem of this section generalizing the results of (Beyer et al., 1999) and (Durrant and Kabán, 2009) to the multi-vector setting.
@@
For sufficiently large databases, a multi-vector search instance is stable if and only if
@@
We defer the proof of this theorem to Appendix C. With this extension of stability to the multi-vector setting in hand, we will now define a procedure for constructing a single-vector search instance from a multi-vector problem.
@@ 4.2 Theoretical Results
A near-neighbor search problem $(Q,D,\delta)$ is $c$-strongly stable for some constant $c>1$ if $\frac{\mathrm{DMAX}}{\mathrm{DMIN}}>c+\Omega(1)$ almost surely.
Intuitively, strong stability asserts that a near neighbor problem is not only stable but also stable with a non-decreasing gap. We will need this stronger property when analyzing the stability of multi-vector search because, even if the induced single-vector search problem is stable, the organization of document vectors into sets might contract the maximum distance gap seen in the single vector search problem. Strong stability ensures that, as long as the vector sets obey a certain mild non-degeneracy condition, we retain stability even if the stability gap contracts.
@@
We provide the proofs of Lemma 4.8 and Theorem 4.9 in Appendix C. In Appendix C.2, we also verify that all three theorem assumptions comfortably hold true in practice.